If Ø(x, y, z) = x² - y²-2²-2, then ∇Ø at (2,1,-1) is ,,,choose the correct answer?


A) 4i +2j-2k

B)4i-2j-2k

C) 4i + 2j + 2k

D)4i - 2j +2k


1
Expert's answer
2021-08-31T23:57:14-0400

By definition, "curl(x\u00b2 - y\u00b2 -z\u00b2-2)=\u2207\u00d7(x\u00b2 - y\u00b2-z\u00b2-2)" or, equivalently,

"∇×(x² - y²-z²-2)= \begin{vmatrix}
 i & j & k\\
 \frac{\partial }{\partial x} & \frac{\partial }{\partial y} & \frac{\partial }{\partial z}\\
x^2 & y^2 & z^2-2
\end{vmatrix}"

"curl(x^2,y^2,z^2-2) =(\frac{∂}{∂y}(y^2)−\frac{∂}{∂z}(z^2-2),\frac{∂}{∂z}(z^2-2)−\frac{∂}{∂x}(x^2),\frac{∂}{∂x}(x^2)−\frac{∂}{∂y}(y^2))\\"

Now, just plug in the found partial derivatives to get the curl:

"curl(x^2,y^2,z^2-2)=(2y-2z,2z-2x,2x-2y)."

Finally, find the curl at the specific point.

"(curl(x^2,y^2,z^2-2))|((x_0,y_0,z_0)=(2,1,-1))=(4,-6,2)"

The answer should be 4i - 6j+2k